# Speculative Model Risk in Healthcare AI: Using Storytelling to Surface Unintended Harms

### v1

@@ 5 User Study
Speculative Model Card Task. This study used a between-subjects design MacKenzie and Castellucci (2016). Each participant completed a speculative model card, a structured template for describing an AI system’s intended use, potential benefits, and possible harms, under one of two conditions. In the Control condition, participants completed the model card directly without storytelling or discussion. In the Story condition, participants viewed the same model card but first joined a red-team discussion on our red-team discussion platform to explore possible benefit and harm stories related to the AI system before completing the documentation. See Figure 9 in the appendix for the model card template.
@@
User Study Results. We conducted a user study with 18 participants to examine how storytelling-based discussions influence ethical reasoning in AI documentation. Participants completed a speculative model card task under two conditions: a control group that worked individually, and a treatment group that used our Story-Driven Red-Team Discussion Room. The discussion platform enabled participants to engage with simulated expert personas in guided, story-based conversations about the potential benefits and harms of AI systems. Each session included three stages: a pre-survey, the model card completion task, and a post-survey evaluating perceived usefulness, trust, and engagement. We analyzed participants’ model card responses (benefit and harm use cases) and post-survey feedback to assess how narrative interaction supported ethical reflection. Results are organized into three key areas: (1) identifying potential benefits, (2) uncovering possible harms, and (3) linking harms to participants’ personal needs and contexts. We applied qualitative coding to classify harm and benefit types, with two annotators achieving moderate agreement (Cohen’s $\kappa=0.460$ for harms and $\kappa=0.476$ for benefits). Study design and full procedures are provided in Appendix A.4.
@@
Does Storytelling Help Identify More Harms? We analyzed responses across 13 harm subtypes defined by Shelby et al. (2023). See Table 7 in the appendix for the full category list. As shown in Table 8, the Control group concentrated on a few categories, mainly privacy violations (37.5%) and diminished health or well-being (20.8%), which together accounted for more than half of all harms. In contrast, the Story condition exhibited a more balanced distribution, with all 13 subtypes represented and none exceeding 16%. Categories such as alienation, stereotyping, and tech-facilitated violence appeared only in the Story group, suggesting that narrative prompts encouraged recognition of subtler and context-dependent harms.
We quantified this difference using Shannon entropy ($H$), which measures distributional diversity across harm types. As shown in Appendix Table 8, entropy increased from 2.433 in the Control condition to 3.383 in Story, and a bootstrap $t$-test confirmed this difference was significant (t=274.15t=-274.15, $p<.001$). These results suggest that storytelling-based discussions broadened participants’ awareness of potential harms and fostered more diverse ethical reasoning.
@@
Does Storytelling Help Reveal More Benefits? We examined whether storytelling broadened participants’ recognition of potential benefits across 12 predefined subtypes, summarized from consumer health AI research Pedroso and Khera (2025); Chustecki (2024). Detailed category descriptions are in Table 9 in Appendix. As shown in Table 10, participants in the Control group focused on a narrow set of advantages, primarily continuous monitoring & self-care (26.1%) and data synergy & learning (21.7%), which together accounted for nearly half of all responses. In contrast, the Story group described a more balanced range of benefits, with all 12 subtypes represented and none exceeding 17%. New themes such as communication & language support, mental health & emotional support, and democratized care & telehealth appeared only under the storytelling condition, suggesting that narrative prompts helped participants consider a wider spectrum of potential advantages.
@@
We quantified this difference using Shannon entropy ($H$), which captures the evenness of the benefit distribution. Entropy increased from 2.579 in the Control condition to 3.554 in the Story condition, and a bootstrap $t$-test confirmed this difference was statistically significant (t=280.87t=-280.87, $p<.001$) in Appendix Table 10. These results indicate that storytelling-based discussions encouraged participants to recognize a more diverse and balanced set of AI benefits.
@@
Does Human–AI Discussion Facilitate Ethical Reflection Through Think-Aloud Engagement? Post-survey responses indicated that Human–AI storytelling discussions fostered deeper ethical and contextual reflection on AI systems. Participants reported that the narrative format helped them articulate risks that were otherwise difficult to express. For example, P9 shared that “It helped me to understand more,” and P7 noted that “The story provides a concrete example of how AI can be harmful.” Engaging with concrete, narrative scenarios prompted participants to articulate their reasoning about model risks in a think-aloud manner, supporting ethical reflection without requiring prior expertise. As P4 explained, “I could not think of [risks] really, but the story shifted my focus to the negative aspect of things which we usually ignore.” Others observed that the stories surfaced overlooked issues, such as “the lack of cultural context” (P6) or emotional harms like “masking of feelings” (P3), suggesting that narrative prompts helped participants identify subtle sociotechnical risks often absent from formal documentation.
@@
Finally, participants found the storytelling approach both engaging and accessible. By embedding risk exploration within narrative contexts, the format allowed learners to focus on ethical reflection rather than technical complexity. As P12 remarked, “It makes the model more interesting and understandable,” and P8 noted that the story “helped me to know how to use the AI tool,” indicating that minimal prior expertise was required to engage meaningfully with ethical scenarios. These findings suggest that Human–AI storytelling discussions can sustain interest while supporting active, reflective engagement with model risks and benefit.
@@ 6 Conclusion
Acknowledgements
This material is based upon work supported by the National Science Foundation (NSF) under Grant No. 2145357.
@@ 7 Limitation
This work has several limitations that indicate directions for future extension rather than weaknesses. Our scenarios focus on consumer health and do not include regulated domains such as clinical decision support, finance, or law. While the framework could be applied to these areas, we have not yet tested it there. The scenarios are synthetic and derived from AI concepts with assistance from LLMs. This enables early exploration of ethical issues but cannot substitute for analysis of deployed systems.
@@ A.2 Case Study
A.3 Experimental Setup and Evaluation
@@
A.4 User Study Design and Procedure
Demographic Attribute Sample (N=18)
@@ A.4 User Study Design and Procedure
   Hispanic or Latino 27.8%
   Asian 44.4%
   Black or African descent 11.1%
   Arab 11.1%
   White 5.6%
@@
Table 5: Demographics of study sample (N=18)
Participants. We recruited 18 participants through university mailing lists and community forums, following screening criteria to ensure informed and reflective discussion. Each participant received a \$10 gift card as compensation for their time. Eligible participants demonstrated prior interest or coursework in Model Cards and Ethical AI. The sample included 11 male and 7 female participants, aged 18–59 ($M=32.6$, $SD=9.9$), representing diverse ethnic backgrounds (Asian: 44.4%, Hispanic/Latino: 27.8%, Black or African descent: 11.1%, Arab: 11.1%, White: 5.6%). See Table 5 for a summary of participant demographics and Table 6 for the distribution across the control and story conditions. Participants were students or professionals in fields such as Computer Science, Data Analytics, Applied Statistics, and Artificial Intelligence. They were randomly assigned to either the story condition ($n=9$) or the control condition ($n=9$). The study was conducted in a hybrid format, with participants joining the Red-Team Discussion Room via computer and interacting with moderators either in person or over Zoom. Survey instruments are detailed in Figure 10 and 11.
@@
P1 Control Female 18-29 Arab Information Technology
P2 Control Female 30-39 Black or African descent Computer science
P3 Story Female 18-29 Asian Information Technology
P4 Control Male 30-39 Hispanic or Latino Data Analytics
P5 Control Female 18-29 Black or African descent Computer science
P6 Control Male 18-29 Arab Masters of Science in Data Analytics
P7 Story Male 18-29 Asian Data Analytics
P8 Story Female 18-29 Asian Data Analytics
P9 Story Female 40-49 Asian Information Technology
P10 Control Male 50-59 Asian Information Technology
P11 Control Male 30-39 Hispanic or Latino Computer science
P12 Control Male 40-49 Hispanic or Latino Information Technology
P13 Story Male 30-39 White Computer science
P14 Story Male 18-29 Hispanic or Latino Computer science
P15 Story Male 30-39 Hispanic or Latino Computer science
P16 Story Male 18-29 Asian Data Analytics
P17 Story Male 30-39 Asian Computer science
P18 Control Female 18-29 Asian Data Analytics
Table 6: Participant demographics by study condition (N=18)
@@
Red-Team Discussion Room Design. The Story-Driven Red-Team Discussion Room is a multi-agent conversational system built on the Cinema of Thought framework Ryu et al. (2025), where users interact with LLM-based agents that embody different personas with distinct ethical perspectives and domain expertise Ryu et al. (2025). Recruiting large, diverse expert groups for red-teaming is costly and logistically challenging. Instead, we simulate expert interactions using multi-agent conversations (GPT-4o-mini) to provide a scalable and accessible alternative. The system combines storytelling, guided prompts, and structured discussions to support ethical reflection and help users explore the consequences of AI behavior from multiple perspectives. Screenshots of the interface are shown in Figure 7 and  8. The corresponding code and prompt can be found in the project’s GitHub repository.
@@
To further support engagement, we provided users with optional hints, short opinion prompts (e.g., “I think…”), follow-up questions (e.g., “Tell me more about…”), and “what if” scenarios to surface potential risks such as bias, misuse, or contextual mismatches. Prior research shows that role-play and narrative methods foster empathy and critical thinking by encouraging users to consider other perspectives Zhang et al. (2025); Ryu et al. (2025). By embedding low-stakes role-play and open-ended ethical questions (e.g., “What could go wrong?” or “Which settings amplify risk?”), the system helps users reflect on how AI behavior varies by context, user, and environment Klassen and Fiesler (2022). Rather than leading users to predefined conclusions, the system encourages them to form their own views, supporting ethical awareness and personal coping strategies through storytelling and simulation.
@@
Experimental Procedure (User Study). In our study, we first conducted a 10-minute orientation to help participants become familiar with the Discussion Room interface, understand the concept of Model Cards and sociotechnical harms, and view a brief demo of a human-AI panel discussion. In the demo, a Medical Expert agent discussed a user story from their professional perspective, and a second expert agent, such as a Research Scientist, Clinical Nurse, or AI Engineer, contributed additional viewpoints. A Moderator agent directed the flow of conversation by posing ethical questions, shifting topics when necessary, and providing reflective prompts. Participants were asked to engage in the discussion by responding to questions or posing their own.
Following the orientation, participants joined in our storytelling‑driven “Red Team Discussion Room” in which each participant was randomly assigned one of three speculative AI model cards: FaceVitals (infers heart rate, blood pressure, and stress from facial video), SensiAI (always on audio and sensor monitoring for older adults with dementia), or CarbLens (estimates carbohydrate intake from meal photos using blood‑glucose and insulin data). Each card was presented with a good user story showing potential benefits and a bad user story highlighting possible harms. These stories served as starting points for discussion, helping participants think critically about potential misuse scenarios, reflect on sociotechnical trade-offs, and articulate ethical concerns related to the selected AI model. In the 20-minute discussion, Participants were asked to engage as they would in a real group discussion, such as responding to others, posing questions, expressing opinions, and collaboratively exploring the issues raised. They were free to ask any questions or respond at any time, with the understanding that there were no right or wrong answers. Participants were encouraged to share any ideas or concerns before the end of the discussion simulation.
And then, participants were instructed to completed a pre-study survey that gathered background information, including demographics, familiarity with AI and model cards, and attitudes toward using stories. They then received a brief overview of the study tasks. In this phase, participants were asked to complete the ethical considerations section of a model card. Specifically, they were required to write at least two “good” and two “bad” use cases based on given a description of a speculative AI system and its intended use. Each use case needs to describe who uses the system, what input it receives, what the AI does, and what the outcome is, highlighting either what goes well or what goes wrong. For each “bad” case, participants were also asked to brainstorm possible mitigation strategies. They were encouraged to think aloud and go beyond the given stories and minimum requirements by generating additional use cases if possible.
After completing the model card task, participants completed Likert-scale survey with open-ended questions to evaluate our storytelling-driven discussion platform. The survey assessed the perceived effectiveness, trustworthiness, satisfaction, and helpfulness of the storytelling-driven discussion platform in supporting model card completion, as well as in brainstorming future use cases and anticipating uncertainties, following established methodological guidelines Kuang et al. (2023). Participants were also asked to reflect on their experience by answering questions about difficult sections of the model card, unclear risks, perceived drivers of AI harms, and how the narrative discussion influenced their understanding or revealed overlooked scenarios. We also collected feedback on system improvements. Participants were asked what features would better support risk brainstorming and how the tool could integrate more effectively with existing documentation workflows. A screenshot of the model card study interface is shown in Figure 9. Transcripts were analyzed using an inductive thematic analysis approach Thomas (2006).
A.5 Additional User Study Findings

### v3

@@ 5 User Study
Speculative Model Card Task. This study used a between-subjects design MacKenzie and Castellucci (2016). Participants completed a speculative model card, a structured template describing an AI system’s intended use, benefits, and potential harms, under one of three conditions. In the Control condition, participants completed the model card directly. In the Story-Only condition, participants first read text-based benefit and harm stories, without discussion, before completing the model card. In the Story condition, participants engaged in a red-team discussion on our platform to explore benefit and harm stories before completing the same model card. The model card template is shown in Figure 7 in the appendix.
@@
User Study Results. We conducted a user study with 45 participants to examine how storytelling-based discussions influence ethical reasoning in AI documentation. Participants completed a speculative model card task under three conditions: a control group that worked individually, a story-only group that read the text-based stories without discussion, and a treatment group that used our Story-Driven Red-Team Discussion Room. The discussion platform enabled participants to engage with simulated expert personas in guided, story-based conversations about the potential benefits and harms of AI systems. Each session included three stages: a pre-survey, the model card completion task, and a post-survey evaluating perceived usefulness, trust, and engagement. We analyzed participants’ model card responses (benefit and harm use cases) and post-survey feedback to assess how narrative interaction supported ethical reflection. As an exploratory qualitative study, we focus on recurring themes rather than statistical power, and prior work shows that small samples are sufficient to reach thematic saturation, where few new themes emerge with additional data Hennink and Kaiser (2022). Results are organized into three key areas: (1) identifying potential benefits, (2) uncovering possible harms, and (3) linking harms to participants’ personal needs and contexts. We applied qualitative coding to classify harm and benefit types, with two annotators achieving moderate agreement (Cohen’s $\kappa=0.4368$ for harms and $\kappa=0.3968$ for benefits). Study design and full procedures are provided in Appendix A.5. As a robustness check, we conducted an LLM-based simulated survey under the same conditions, as described in Appendix A.7.
@@
Does Storytelling Help Identify More Harms? We analyzed responses across 17 harm subtypes defined by Shelby et al. (2023). See Table 7 in the appendix for the full category list. As shown in Table 8, the Control group concentrated on a small set of categories, primarily diminished health or well-being (32.3%), service or benefit loss (24.2%), and privacy violations (22.6%), which together accounted for the majority of reported harms. In contrast, both the Story-only and Story conditions exhibited more distributed coverage across harm types, with the Story condition showing the widest range of subtypes and lower concentration in any single category. Several harm categories, including cultural harms, political and civic harms, and tech-facilitated violence, appeared only in the Story condition, suggesting that interactive narrative discussion supported recognition of less obvious and context-dependent harms. We quantified these differences using Shannon entropy ($H$), which measures distributional diversity across harm types. As shown in Table 8, entropy increased from 2.329 in the Control condition to 2.927 in the Story-only condition and to 3.701 in the Story condition. Bootstrap $t$-tests confirmed that both Story-only and Story exhibited significantly higher entropy than Control ($p<.001$), and that Story also showed significantly higher entropy than Story-only ($p<.001$). These results indicate that storytelling-based engagement, particularly interactive discussion, broadened participants’ awareness of potential harms and supported more diverse ethical reasoning.
@@
Does Storytelling Help Reveal More Benefits? We examined whether storytelling broadened participants’ recognition of potential benefits across 18 predefined subtypes, summarized from prior consumer health AI research Pedroso and Khera (2025); Chustecki (2024). Detailed category descriptions are provided in Table 9 in the appendix. As shown in Table 10, participants in the Control group concentrated on a small set of benefits, primarily decision support & diagnostic augmentation (25.4%), continuous monitoring & self-care (23.8%), and early detection & prediction (22.2%), which together accounted for the majority of responses. In contrast, both the Story-only and Story conditions exhibited broader coverage across benefit types, with the Story condition showing the most diverse distribution and lower concentration in any single category. Several benefit subtypes, including accessibility & disability support, clinician workload relief, and transparency & trust, appeared only in the Story condition, suggesting that interactive narrative discussion supported recognition of less salient or less frequently considered benefits.
@@
We quantified these differences using Shannon entropy ($H$), which captures the evenness of the benefit distribution. As shown in Table 10, entropy increased from 2.407 in the Control condition to 3.242 in the Story-only condition and to 3.868 in the Story condition. Bootstrap $t$-tests confirmed that both Story-only and Story exhibited significantly higher entropy than Control ($p<.001$), and that Story also showed significantly higher entropy than Story-only ($p<.001$). These results indicate that storytelling-based engagement—particularly interactive discussion—encouraged participants to recognize a more diverse and balanced set of potential AI benefits.
@@
What do People Say About Ethical Reflection in Speculative AI Documentation? Post-survey responses indicated that Human–AI storytelling discussions fostered deeper ethical and contextual reflection on AI systems. Participants reported the narrative format helped them articulate risks that were otherwise difficult to express. For example, P9 shared that “It helped me to understand more,” and P7 noted that “The story provides a concrete example of how AI can be harmful.” Engaging with concrete narrative scenarios led participants to verbalize their reasoning about model risks in a think-aloud manner, supporting ethical reflection without requiring prior expertise. As P4 explained, “I could not think of [risks] really, but the story shifted my focus to the negative aspect of things which we usually ignore.” Others observed that stories surfaced overlooked issues, such as “the lack of cultural context” (P6) or emotional harms like “masking of feelings” (P3), suggesting that narrative prompts helped surface subtle sociotechnical risks often missing from formal documentation.
@@
Finally, participants found the storytelling approach both engaging and accessible. By embedding risk exploration within narrative contexts, the format allowed learners to focus on ethical reflection rather than technical complexity. As P12 remarked, “It makes the model more interesting and understandable,” and P8 noted that the story “helped me to know how to use the AI tool,” indicating that minimal prior expertise was required to engage meaningfully with ethical scenarios. These findings suggest that Human-AI storytelling discussions can sustain interest while supporting active, reflective engagement with model risks and benefit.
@@ 7 Limitation
This work has several limitations that indicate directions for future extension rather than weaknesses. Our scenarios focus on consumer health and do not include regulated domains such as clinical decision support, finance, or law. While the framework could be applied to these areas, we have not yet tested it there. The scenarios are synthetic and derived from AI concepts with assistance from LLMs, enabling early exploration of ethical issues but not substituting for analysis of deployed systems. We intentionally use low-stakes, synthetic consumer-health scenarios and university participants to validate the storytelling method before deployment in regulated or clinical environments. This study should therefore be understood as a first step that demonstrates methodological feasibility rather than direct applicability to clinicians or patients, which we leave to future work.
@@ A.2 Case Study
A.3 Alignment Between Human and LLM-Based Evaluations
Our study compares human preferences with LLM-as-a-judge at the overall method level. As reported in Section 4, human annotators strongly prefer our Storytelling method (88% for Llama3 and 76% for Gemma). This preference closely aligns with results obtained using GPT-4o as an automatic judge. Both humans and LLM judges reach the same conclusion: our method produces higher-quality stories than the baseline and ablation variants. We also observe consistent stylistic patterns across evaluators. Human annotators slightly prefer Llama3 because its stories are easier to follow, whereas Gemma tends to generate more expressive and detail-dense narratives, a distinction that is also reflected in GPT-4o’s scores.
A more fine-grained comparison between human and LLM judges at the level of individual dimensions (e.g., creativity, coherence, engagement) could provide additional insight. We did not collect dimension-specific human ratings because doing so would impose substantial cognitive load on annotators in this early-stage study, which was designed to focus on overall story quality. We will clarify this design choice in the revision and identify dimension-level human–LLM agreement as an important direction for future work. Overall, our results show strong alignment between human judgments and LLM judges in terms of global ranking and observed stylistic tendencies.
Below, we provide a qualitative illustration that helps explain the small differences between human and LLM preferences.
Coherent and Easy to Follow (Typical of Llama3).
Dr. Rivera used DeepHeart to review Ms. Chen’s annual check-up. The system flagged her as “High Risk for Heart Disease” due to low step count and elevated resting heart rate. Ms. Chen explained she stays active through daily childcare and household tasks. Dr. Rivera incorporated this context and updated the assessment.
This narrative is concise and highlights a clear causal chain, making it easy for human annotators to read and interpret.
Detail-Dense and Expressive (Typical of Gemma).
Dr. Ramirez used DeepHeart to assess Miguel Garcia’s metabolic risk. The system labeled him “High Risk for Type 2 Diabetes,” citing low ambulatory activity, irregular heart-rate variability, and disrupted sleep cycles. Miguel described demanding overnight construction work, inconsistent shift schedules, and traditional dietary practices that the model misinterpreted. These omissions led to unnecessary tests and referrals, leaving Miguel frustrated.
This narrative contains substantially more physiological, contextual, and cultural detail. LLM judges often reward this richness, whereas human annotators sometimes find such stories harder to follow.
In summary, humans slightly prefer Llama3 due to its clarity and readability, while LLM-as-a-judge occasionally favors Gemma for its more elaborate and detail-dense narratives. Despite these differences, both humans and LLM judges agree on the central result: the Storytelling method performs best overall.
A.4 Experimental Setup and Evaluation
@@
A.5 User Study Design and Procedure
Demographic Attribute Sample (N=45)
@@ A.5 User Study Design and Procedure
   Hispanic or Latino 20.0%
   Asian 44.5%
   Black or African descent 6.7%
   Arab 4.4%
   White 24.4%
@@
Table 5: Demographics of study sample (N=45)
Participants. We recruited 45 participants through university mailing lists and community forums, following screening criteria to ensure informed and reflective discussion. Each participant received a \$10 gift card as compensation for their time. Eligible participants demonstrated prior interest or coursework in Model Cards and Ethical AI. The sample included participants of diverse genders (66.7% male, 33.3% female), ages ranging from 18 to 60+ (with the majority between 18–39), and diverse ethnic backgrounds (Asian: 44.5%, Hispanic or Latino: 20.0%, White: 24.4%, Black or African descent: 6.7%, Arab: 4.4%). See Table 5 for a summary of participant demographics and Table 6 for the distribution across the three study conditions. Participants were students or professionals in fields such as Computer Science, Data Analytics, Applied Statistics, and Artificial Intelligence. Participants were randomly assigned to one of three conditions: a control condition, a story-only condition, or a story-driven discussion condition. The study was conducted in a hybrid format, with participants joining the Red-Team Discussion Room via computer and interacting with AI moderators either in person or over Zoom. Survey instruments are detailed in Figure 10 and Figure 11.
@@
P1 Control Female 18-29 White Music
P2 Control Male 40-49 Black or African descent Electrical Engineering
P3 Control Female 18-29 Arab Information Technology
P4 Control Female 40-49 White Nursing
P5 Control Male 18-29 Asian Information Technology
P6 Control Male 30-39 Asian Computer science
P7 Control Female 18-29 Black or African descent Computer science
P8 Control Male 18-29 Arab Data Analytics
P9 Control Male 30-39 Hispanic or Latino Data Analytics
P10 Control Male 30-39 Hispanic or Latino Computer science
P11 Control Male 40-49 Hispanic or Latino Information Technology
P12 Control Male 18-29 Hispanic or Latino Law
P13 Control Male 50-59 Asian Information Technology
P14 Control Female 18-29 Asian Data Analytics
P15 Control Female 30-39 Black or African descent Computer science
P16 Story_only Female 18-29 Asian Information Technology
P17 Story_only Male 30-39 White Information Technology
P18 Story_only Male 30-39 Asian Computer science
P19 Story_only Male 30-39 Asian Electrical Engineering
P20 Story_only Male 30-39 Asian Computer science
P21 Story_only Female 30-39 Asian Electrical Engineering
P22 Story_only Male 60+ Hispanic or Latino Computer science
P23 Story_only Male 18-29 Asian Information Technology
P24 Story_only Male 18-29 White Computer science
P25 Story_only Female 18-29 Asian Computer science
P26 Story_only Female 18-29 Hispanic or Latino Education
P27 Story_only Male 18-29 White Information Technology
P28 Story_only Male 18-29 White Computer science
P29 Story_only Male 30-39 Asian Information Technology
P30 Story_only Male 30-39 Asian Electrical Engineering
P31 Story Female 60+ White Nursing
P32 Story Male 18-29 White Computer science
P33 Story Male 18-29 Hispanic or Latino Computer science
P34 Story Female 18-29 Asian Information Technology
P35 Story Male 18-29 White Information Technology
P36 Story Male 18-29 Hispanic or Latino Computer science
P37 Story Male 30-39 Asian Computer science
P38 Story Male 30-39 Hispanic or Latino Computer science
P39 Story Female 40-49 Asian Information Technology
P40 Story Female 30-39 Asian Data Analytics
P41 Story Female 18-29 Asian Data Analytics
P42 Story Male 18-29 Asian Data Analytics
P43 Story Male 18-29 Asian Data Analytics
P44 Story Male 60+ White Medicine
P45 Story Male 30-39 White Computer science
Table 6: Participant demographics by study condition (N=45)
Experimental Procedure (User Study). All participants, regardless of condition, began with a standardized introductory tutorial led by a graduate student researcher. The tutorial lasted approximately 15 minutes and introduced the concept of model cards, key ethical and sociotechnical considerations, example benefit and harm use cases, and instructions for completing the model card task. The Control condition received a version consisting of approximately 20 slides, followed by a brief Q&A session to ensure task understanding. The Story and Story-only condition received a slightly extended version (approximately 25 slides), which included five additional slides introducing storytelling as a lens for reasoning about AI harms and benefits. This shared tutorial and Q&A ensured that both groups received comparable guidance, examples, and preparation prior to the main task.
After the tutorial, participants proceeded according to condition. Participants in the Story condition viewed a storytelling-driven “Red Team Discussion Room” simulation. In this approximately 15-minute session, participants observed a speculative human–AI panel discussion centered on a single AI system. A Moderator agent guided the conversation by posing ethical questions, shifting topics as needed, and offering reflective prompts. Two expert agents (e.g., a Medical Expert and a Research Scientist, Clinical Nurse, or AI Engineer) discussed the system from complementary professional perspectives. Participants were encouraged to engage as they would in a real group discussion by responding to questions, expressing opinions, or posing their own questions.
Participants in the Story-Only condition received the same speculative narratives but without the discussion interface or dialogue. Instead, they were shown a static textual presentation consisting of one good and one bad user story describing the same AI system. They did not observe or participate in any panel discussion and did not interact with the story content beyond reading it. Participants in the Control condition did not receive any storytelling component beyond the examples included in the tutorial.
In both Story and Story-Only conditions, each participant was randomly assigned one of three speculative AI model cards: Moodcapture (infers heart rate, blood pressure, and stress from facial video for detect emotion), SensiAI (always-on audio and sensor monitoring for older adults with dementia), or Gluroo Ai (estimates carbohydrate intake from meal photos using blood-glucose and insulin data). Each participant viewed exactly one pair of narratives (one benefit and one harm) associated with a single AI system. Participants did not navigate among multiple stories or interactively explore alternative scenarios. These narratives served as conceptual anchors for reflecting on potential misuse scenarios, sociotechnical trade-offs, and ethical risks.
Following the tutorial and condition-specific materials, all participants completed a pre-study survey collecting background information, including demographics, familiarity with AI and model cards, and attitudes toward using stories in ethical reasoning. Participants then completed the core model card task. Specifically, they were asked to fill out the ethical considerations section of a speculative model card by writing at least two good and two bad use cases. Each use case was required to describe (1) who uses the system, (2) what input it receives, (3) what the AI does, and (4) the resulting outcome, highlighting either positive or negative consequences. For each “bad” use case, participants were also asked to propose possible mitigation strategies. Participants were encouraged to think aloud and to generate additional use cases beyond the minimum requirements if possible.
After completing the model card task, participants filled out a post-study survey consisting of Likert-scale items and open-ended questions assessing the perceived effectiveness, trustworthiness, satisfaction, and helpfulness of the study materials in supporting model card completion, brainstorming future use cases, and anticipating uncertainties, following established methodological guidelines Kuang et al. (2023). Participants also reflected on which sections of the model card they found most challenging, which risks remained unclear, perceived drivers of AI harms, and how (if applicable) the narrative materials influenced their understanding or revealed overlooked scenarios. We additionally collected feedback on desired system improvements and how such tools might better integrate with existing documentation workflows. A screenshot of the model card study interface is shown in Figure 7. All discussion transcripts and open-ended responses were analyzed using an inductive thematic analysis approach Thomas (2006).
Timing. The total session duration was approximately 30–45 minutes for participants in the Control condition, approximately 40–50 minutes for participants in the Story-Only condition, and approximately 45–55 minutes for participants in the Story condition, reflecting the additional discussion component.
@@
Red-Team Discussion Room Design. Participants assigned to the Story condition interacted with the Story-Driven Red-Team Discussion Room, a multi-agent conversational system built on the Cinema of Thought framework Ryu et al. (2025). The system enables participants to engage with LLM-based agents that embody distinct personas with diverse domain expertise and ethical perspectives, supporting structured reflection on potential benefits, harms, and sociotechnical trade-offs of AI systems. Recruiting large, diverse expert groups for red-teaming is costly and logistically challenging. Instead, we simulate expert interactions using multi-agent conversations (GPT-4o-mini) to provide a scalable and accessible alternative. The system combines storytelling, guided prompts, and structured discussions to support ethical reflection and help users explore the consequences of AI behavior from multiple perspectives. Screenshots of the interface are shown in Figure 8 and  9. The corresponding code and prompt can be found in the project’s GitHub repository.
@@
To further support engagement, we provided users with optional hints, short opinion prompts (e.g., “I think…”), follow-up questions (e.g., “Tell me more about…”), and “what if” scenarios to surface potential risks such as bias, misuse, or contextual mismatches. Prior research shows that role-play and narrative methods foster empathy and critical thinking by encouraging users to consider other perspectives Zhang et al. (2025b); Ryu et al. (2025). By embedding low-stakes role-play and open-ended ethical questions (e.g., “What could go wrong?” or “Which settings amplify risk?”), the system helps users reflect on how AI behavior varies by context, user, and environment Klassen and Fiesler (2022). Rather than leading users to predefined conclusions, the system encourages them to form their own views, supporting ethical awareness and personal coping strategies through storytelling and simulation.
@@
A.6 Additional User Study Findings